OpenStudy (anonymous):

(intergral from 1 to infinity) [ln(x)]/[x^2] convergent or divergent?

OpenStudy (anonymous):

i would say converge, but lets check it

OpenStudy (anonymous):

I know it converges I just don't know how to show it does

OpenStudy (anonymous):

find the anti derivative think parts is what you need

OpenStudy (anonymous):

i did parts and i got (-ln(x)/x) from one to infinity -1/x from one to infinity

OpenStudy (anonymous):

put it over one denominator to make your life easier then take the limit as x goes to infinity and see you get zero

OpenStudy (anonymous):

so it would be (-ln(x)-1)/x

OpenStudy (anonymous):

\[\lim_{x\to \infty}-\frac{\ln(x)+1}{x}=0\] for sure

OpenStudy (anonymous):

why is the one not negative?

OpenStudy (anonymous):

i pulled the minus sign out all the way no difference

OpenStudy (anonymous):

ok so we get negative infinity over infinity

OpenStudy (anonymous):

ok now i suppose if you want to make your teacher happy you could appeal to l'hopital's rule, though it should be clear that \(\ln(x)\) grows much much more slowly than \(x\)

OpenStudy (anonymous):

so it would be -1/x from 1 to infinity?

OpenStudy (anonymous):

the limit will be zero you still have to compute the integral, i.e. plug in 1

OpenStudy (anonymous):

I don't understand. if we use l'hopital's rule to get -1/x from one to infinity i'm getting -1 as an answer

OpenStudy (anonymous):

hold on i think i have confused you, let's take it step by step

OpenStudy (anonymous):

oh nevermind I got it -1/ infinity goes to zero then minus -1/1 haha brain fart

OpenStudy (anonymous):

ok good got it

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

is there a way I can thank you on here except for just saying thank you?